In a triangle, the measure of the first angle is 15degrees more than twice the measure of the second angle. The measure of the third angle exceeds that of the second angle by 25 degree. What is the measure of each angle ?
Answer by Boreal(15235)

second angle is x
first angle is 2x+15
third angle is x+25
One can make any angle x, but I like to make the smallest x and then work from there.
These add to 180 degrees.
x+2x+15+x+25=180
4x+40=180
4x=140
x=35 degrees, angle 2
2x+15=85 degrees, angle 1
x+25=60 degrees, angle 3
They add to 180 degrees.